# Deep Time Series Models: A Comprehensive Survey and Benchmark

### v1

@@ 5.2 Evaluation Protocols
For long-term forecasting and imputations, we rely on Mean Square Error (MSE) and Mean Absolute Error (MAE) as the primary evaluation metrics. These metrics help us accurately assess the accuracy of our predictions and imputations. For short-term forecasting, we use the Symmetric Mean Absolute Percentage Error (SMAPE) and Mean Absolute Scaled Error (MASE) as metrics, which focus on absolute errors and reduce the impact of outliers, providing reliable evaluations of forecast accuracy across different datasets and methodologies. In the case of time series classification tasks, we utilize Accuracy as the evaluation metric. Accuracy measures the overall prediction performance by calculating the ratio of correctly classified samples to the total number of samples. For anomaly detection, we employ the F1 score to validate the identification of abnormal values. The F1 score represents a balanced combination of precision and recall, offering a comprehensive assessment of a classifier’s performance, especially when dealing with imbalanced classes in the context of anomaly detection.
5.3 Datasets
TSLib includes a variety of mainstream datasets across different numbers of samples and categories, a richness of tasks, and a diversity of domains. In this section, we will focus on introducing representative datasets from various time series analysis tasks included in the Time Series Library.
Classification Time series classification aims to assign a label or category to a time series based on its temporal features. To evaluate this capability, we selected ten multivariate datasets from the UEA Time Series Classification Archive [164], supported in our TSLib. These datasets cover a range of practical tasks, including gesture, action, audio recognition, and medical diagnosis through heartbeat monitoring. We pre-processed the datasets according to the descriptions provided in [165]. Detailed dataset descriptions are shown in Table III.
Imputation Due to glitches, the collected time series data may contain partially missing values, posing a challenge for time series analysis. Therefore, time series imputation is a crucial task in real-world applications, aiming to fill in missing values within a time series based on contextual observations from the data. For our benchmark, we selected Electricity Transformer Temperature (ETT) [21], Electricity, and Weather to evaluate the performance of time series imputation tasks with different missing ratios.
@@ 5.3 Datasets
Forecasting Time series forecasting is an essential task in time series analysis, and it has been widely explored in academic and industry domains. By leveraging historical patterns and trends, the model can predict future values or trends of a time series. Time series forecasting can be broadly divided into two types: long-term forecasting and short-term forecasting. For the long-term time series forecasting task, a wide range of datasets are included in our benchmark, including Electricity Transformer Temperature (ETT), Electricity, Weather, Traffic, Exchange, and Illness (ILI). For the short-term forecasting task, we selected the M4 dataset [166], which comprises six sub-datasets with varying sampling frequencies and domains.
@@
5.4 Main Results
@@
To examine the strengths and limitations of various methods in mainstream time series analysis tasks, we select twelve representative models from our TSLib. These models encompass four popular deep model architectures and address tasks, including long-term and short-term forecasting, classification, imputation, and anomaly detection.
@@
Baselines To conduct a fair comparative analysis and thoroughly explore the effectiveness of different model architectures in various time series analysis tasks, we conduct comparative comparison experiments using state-of-the-art models designed based on different deep architectures. As shown in Figure 11, we select several advanced and representative Transformer-based models: iTransformer [30], PatchTST [23], Autoformer [22], Non-Stationary Transformer (Stationary) [41], and FEDformer [86] to verify the performance. Additionally, we consider TimesNet [3] and SCINet [128] as the CNN-based models to compare. For RNN-based models, we included the novel and effective Mamba [120]. Finally, we include DLinear [52], N-BEATS [24], and TiDE [97] as representative MLP-based models for analysis. It is important to mention that TiDE [97] is designed to be dependent on specific timestamps, and cannot be easily adapted to part tasks without timestamps, such as short-term forecasting, anomaly detection, and classification.
Unified Experimental Settings For the long-term forecasting task, we conducted two experimental settings to ensure a fair and comprehensive comparison: unified hyperparameter and hyperparameter searching. For the unified hyperparameter, we conducted experiments using a range of standardized hyperparameters set across different datasets. This allows us to accurately evaluate the relative performance of time series models with different deep architectures while keeping other factors constant. As for the “hyperparameter searching” scenario, we conducted separate hyperparameter searches for different model architectures and time series datasets. This approach enabled us to identify the best performance of different time series analysis models. By employing above both settings, we obtain a comprehensive understanding of the forecasting performance of different time series models. For the remaining tasks, we maintained the standard experimental settings as outlined in [3] to validate the performance of different time series models.
Overall Results Based on the overall results of the models across different architectures in Figure 11, we surprisingly find that the MLP-based models, which are generally simpler and have lower computational overhead, perform well on the time series forecasting task. However, these models appear to be less effective in other types of tasks, which requires the model to learn more informative representations. On the contrary, the CNN-based models demonstrate more comprehensive capabilities and excel in classification, imputation, and anomaly detection tasks. The RNN-based models, while performing well on anomaly detection tasks, show limited effectiveness compared to other model architectures. In contrast, the Transformer-based models have demonstrated highly competitive performance across various time series analysis tasks. This can be attributed to the powerful data modeling capabilities inherent in the transformer architecture, which contribute to its overall and consistently superior performance across diverse time series analysis tasks. It further shows that Transformer-based models hold significant research value and application potential in the field of time series analysis and have emerged as a particularly promising option in the time series domain.
@@
As illustrated in Figure 10, we have also included more detailed results and a top three performance leaderboard for four representative time series analysis tasks. These results clearly show that the Transformer-based models, namely iTransformer [30] and PatchTST [23], exhibit superior forecasting capabilities compared to other models for both long-term and short-term forecasting tasks. This further proves that it is of great significance and value to explore different modeling methods of temporal tokens in time series. Additionally, TimesNet [3] shows a more comprehensive and effective performance covering time series classification, imputation, and anomaly detection tasks. It has pioneered a milestone in the general time series analysis model.
@@
We believe that TSLib can provide useful start code, valuable insights on model properties and model selection guidance for future research and real-world applications.
@@ 6 Future Directions
In this section, we provide a discussion on the promising directions for time series analysis.
@@ 6.1 Time Series Pre-training
A pretraining-finetuning learning paradigm is a two-stage approach commonly used in nature language processing (NLP) [171], [16], [172], [173], [173] and computer vision (CV) [174], [175], [176], [177]. Pre-training establishes the basis of the abilities of Large Models through unsupervised learning [178, 179]. Fine-tuning can improve the performance of the pre-trained model on a specific task or domain.
@@
Due to the limited availability of labeled datasets, self-supervised pre-training [175] has garnered significant attention and has been extensively investigated in the domains of natural language modeling and computer vision. Self-supervised pre-training paradigm significantly reduces labeling expenses and benefits for diverse downstream tasks. Notably, recent research efforts have introduced several self-supervised pre-training methods tailored for time series data, which can be primarily classified into contrastive learning [180] and masked time series modeling [16, 181].
@@
Contrastive learning refers to learning the representations of data by contrasting between similar and dissimilar pairs, where similar sample pairs are learned to be close to each other and dissimilar pairs are far apart [180, 182, 183]. Although SimCLR [182] has demonstrated remarkable success in the domain of computer vision, directly applying SimCLR to the field of time series data yields unsatisfactory results due to the insufficient modeling of temporal dependencies. CPC [184] introduced contrastive predictive coding, which utilizes model-predicted features as positive samples and randomly-sampled features as negative samples, to obtain time series representations that are advantageous for downstream tasks. TimeCLR [185] proposes a DTW data augmentation to generate phase shift and amplitude-change phenomena which can preserve time series structure and feature information. TS-TCC [186] employs efficient data augmentations designed for time-series data, and learns discriminative representations from the proposed Temporal Contrasting module and Contextual Contrasting module. TS2Vec [187] employs a hierarchical contrastive learning method and defines the contrastive loss from both instance-wise and patch-wise perspectives across different augmented context views, resulting in a robust contextual representation for each timestamp. Furthermore. LaST [188] takes the idea of variational inference theory [189] and proposes seasonal-trend representations learning and disentanglement mechanisms. CoST [190] proposes a contrastive learning framework to learn disentangled seasonal-trend representations for long sequence time series data. TF-C [191] develop frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency.
@@
Masked modeling is a reconstruction-based method, which can predict a masked token in a sequence based on the context unmasked part [192, 16, 181]. TST [193] follows the pre-training paradigm proposed in BERT [16] and proposes a pre-training framework for multivariate time series. Further, PatchTST [23] segments the time series data into multiple non-overlapping patches and proposes a patch-level masked modeling approach. SimMTM [194] relates masked modeling to manifold learning and presents a neighborhood aggregation design for reconstruction based on the similarities learned in series-wise representation space. HiMTM [195] proposes a novel hierarchical masked time series pre-training framework to effectively capture the multi-scale characteristics of time series data. TimeSiam [196] constructs an asymmetric masking reconstruction task to capture intrinsic temporal correlations between randomly sampled past and current subseries and learn internal time-dependent representations based on Siamese networks.
@@ 6.2.1 Time Series Foundation Models
Recent advancements in deep learning, particularly with the emergence of foundation models (FMs), have demonstrated significant progress in natural language processing (NLP) and computer vision (CV) domains [197, 179, 198]. Different from prior deep models, foundation models are pre-trained on massive amounts of data, which enables them to have a wide range of general knowledge learned from diverse domains. Given their success in capturing contextual information and semantic understanding, it is promising to explore a generalized time series foundation model that can effectively learn complex temporal dependencies and capture the underlying dynamics inherent in time series data. Early attempts such as TimeGPT [199], Lag-LlaMa [200], and Timer [155] focus solely on univariate time series data. Nevertheless, in real-world forecasting scenarios, it is crucial to involve additional information that is related to the temporal variation of the target time series and must be taken into account, such as weather conditions or holidays. MOIRAI [201] tries to flatten multivariate time series into a single sequence containing all variate, but its generalization capabilities to other downstream analysis tasks are under-explored. In addition to modeling inter-series dependencies, modeling the relationship between time series and external factors in other can achieve a better understanding of time series data. These external factors may be in the form of other modalities, such as text data or calendar data, and thus multimodal learning is a future trend in the development of a multi-modal time series foundation model.
@@ 6.2.2 Adaptation of Large Language Models
Fine-Tuning Pre-trained Language Models Based on the similar sequential nature, fine-tuning pre-trained language models to equip them with time series analysis capabilities has become a promising research topic. When applying LLMs to time series, it is essential to tokenize the time series before feeding it to a pre-trained model. Thus, adapting an LLM for time series included two key components: time series tokenization and efficient fine-tuning methods for time series analysis tasks. LLM4TS [204] proposes a two-stage fine-tuning strategy, including the time-series alignment stage to align LLMs with the nuances of time series data, and the fine-tuning stage for downstream tasks. LLMTime [205] treats time series forecasting as next-token prediction in text and attempts to encode time series data as a string of numerical digits. Recent Chronos [206] introduces a pre-trained probabilistic time series model based on existing Transformer-based language model architectures. Technologically, Chronos tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the model on these tokenized time series via the cross-entropy loss. Benefiting from the generative capability of LLMs, most of the existing research focuses on time series forecasting tasks. GPT4TS [207] propose a unified framework for diverse time series analysis tasks by using a pre-trained GPT-2 model and fine-tuning the positional embeddings and the parameters of the layer normalization for each analysis task.
Prompting Large Language Models Recent Large Language Models exhibit the strong abilities of in-context learning [146] and instruction following [208]. Therefore, the paradigm of leveraging natural language instructions or task examples to guide the model in addressing novel tasks has emerged as a groundbreaking approach [209, 210, 211], which has become a potential solution for time series analysis tasks [212]. Recent literature, such as PromptCast [212], UniTime [213], and TimeLLM [214] focus on investigating a prompt template to enable LLMs to perform the forecasting task. There are other works represented by Autotimes [55, 215], that attempt to design soft prompts for time series data. However, existing prompting approaches are tailored for forecasting, and how to empower LLMs to other time series tasks besides forecasting is relatively unexplored.

### v2

@@ 5.2 Evaluation Protocols and Metrics
Following previous research, for long-term forecasting and imputations, we adopt Mean Square Error (MSE) and Mean Absolute Error (MAE) as the primary evaluation metrics. For short-term forecasting, we use the Symmetric Mean Absolute Percentage Error (SMAPE) and Mean Absolute Scaled Error (MASE) as metrics, which focus on absolute errors and reduce the impact of outliers, providing reliable evaluations of forecast accuracy across different datasets and methodologies. In time series classification, we utilize classification accuracy as the evaluation metric. For anomaly detection, we employ the F1 score to validate the identification of abnormal values. The F1 score represents a balanced combination of precision and recall, offering a comprehensive assessment of a classifier’s performance, especially when dealing with imbalanced abnormal classes.
5.3 Supported Datasets
TSLib natively supports a variety of widely used datasets across different numbers of samples and variables, covering diverse domains and tasks. The correspondences between tasks and datasets are depicted in Figure 8. Since all datasets are commonly used, here is only a brief introduction, and more details can be found in previous papers [3].
@@ 5.3 Supported Datasets
Forecasting Forecasting is one of the most widely explored tasks in previous research, bringing us a wide choice of experimental datasets. Just as stated in design principles, we further divide this task into two types (long- and short-term) according to the forecasting horizon. For long-term time series forecasting, TSLib includes nine datasets, including Electricity Transformer Temperature (ETT, including four subsets) [21], Electricity, Weather, Traffic, Exchange, and Illness [22]. For the short-term forecasting task, we include the M4 dataset [170], which comprises six sub-datasets with varying sampling frequencies and domains.
Imputation Due to glitches, the collected time series data may contain partially missing values, posing a challenge for time series analysis. However, in real-world applications, it is really hard to obtain the ground truth values for the missing records. Thus, we utilize the Electricity Transformer Temperature (ETT, with four subsets) [21], Electricity, and Weather to construct the missing value situation by randomly masking a certain portion of time steps.
@@
Classification We selected ten multivariate datasets from the UEA Time Series Classification Archive [175] to test the series-level classification tasks. These datasets cover a range of practical tasks, including gesture, action, audio recognition, and medical diagnosis through heartbeat monitoring. We pre-processed the datasets according to the descriptions in [176].
5.4 Default Experimental Setups
Here, we provide experimental setups for each analysis task supported by TSLib. For long-term forecasting, we use a look-back length of 36 and prediction lengths of for the ILI dataset. For other datasets, a unified look-back length of 96 is adopted, with forecast horizons of . In short-term forecasting, the prediction lengths are set as for yearly, quarterly, monthly, weekly, daily, and hourly data, respectively. In the classification task, we aim to categorize high-dimensional time series data into multiple classes. In the imputation task, we randomly mask the time points in ratios of . In the anomaly detection task, the reconstruction is a classical task for unsupervised point-wise representation learning, where the reconstruction error is a natural anomaly criterion. Therefore, we set deep time series models as the base models for reconstruction and utilize the reconstruction error as a shared anomaly criterion across all experiments.
| Task | LR∗ | Loss | BatchSize | Epochs |
| Long-term Forecasting | MSE | 32 | 10 | |
| Short-term Forecasting | SMAPE | 16 | 10 | |
| Imputation | MSE | 16 | 10 | |
| Classification | Cross Entropy | 16 | 100 | |
| Anomaly Detection | MSE | 128 | 10 |
Implementation Details In this paper, all experiments were implemented using PyTorch and executed on an NVIDIA A100 GPU. The Adam optimizer was used for model optimization. Each model’s performance was comprehensively evaluated across various time series analysis tasks. To achieve optimal performance and ensure fair comparisons across models and datasets, we conducted hyperparameter searches for each model architecture following [177]. Specifically, we search the number of hidden layers and the dimension of hidden representations from the range . For each time series analysis task, TSLib utilized task-specific and model-shared training configurations, including learning rate, loss function, batch size, and number of epochs, which are listed in Table IV.
Look-back Searching Setting Note that in long-term forecasting tasks, the model’s performance is highly dependent on the historical length of the input series. Given the inconsistent look-back length of various papers, we further conducted an additional set of experiments referred to as “Look-back Searching” to ensure a fair and reliable comparison. In these experiments, we systematically searched for input length among following [177] and performed an extensive hyperparameter search on each model’s architecture to reach their optimal performance. This approach allowed us to assess the maximum achievable performance of each model while ensuring that the reported results accurately reflect their capabilities and facilitate a fair comparison across different models.
5.5 Model Analysis
@@
To examine the strengths and limitations of various methods in mainstream time series analysis tasks, we select thirteen representative models from TSLib for comparison. These models encompass five popular deep model architectures and tasks, including long-term and short-term forecasting, classification, imputation, and anomaly detection.
@@
Baselines To thoroughly explore the effectiveness of different model architectures in various tasks, we conduct comparative experiments using state-of-the-art models designed based on different deep architectures. As shown in Figure 9, we select several advanced and representative Transformer-based models: iTransformer [39], PatchTST [23], Autoformer [22], Non-Stationary Transformer (Stationary) [38], and FEDformer [83] to verify the performance. Additionally, we consider TimesNet [3] and SCINet [130] as the CNN-based models to compare. For RNN-based models, we included the novel and effective Mamba [118]. As for GNN-based models, since most approaches require a dependency graph as priors for the model, we take StemGNN as the representative model, which enables a data-driven construction of dependency graphs. DLinear [50], N-BEATS [24], and TiDE [96] are tested as representative MLP-based models. Since TiDE [96] depends on timestamps and cannot be easily adapted to tasks without timestamps in their datasets, such as short-term forecasting, anomaly detection, and classification, we only test it on the long-term forecasting task.
@@
Overall Results As shown in Figure 9, we can find that the Transformer-based models, e.g., iTransformer [39] and PatchTST [23], exhibit superior forecasting capabilities compared to other models for both long-term and short-term forecasting tasks. This further proves that it is of great significance to explore time series tokenization strategies, which is the key to unlocking the capacity of Transformer-based models. Additionally, TimesNet [3] shows a more comprehensive and effective performance covering time series classification, imputation, and anomaly detection tasks, pioneering a milestone in task-general time series models.
Beyond detailed comparison for each specific model, we also include a summarized performance comparison among different types of deep learning backbones in the left of Figure 10, where the performance of each type of backbone is the best performance achieved among all corresponding models. We surprisingly find that the MLP-based models, which are generally simpler and have lower computational overhead, perform well on time series forecasting. However, these models appear to be less effective in other types of tasks, which require the model to learn more informative representations. On the contrary, the CNN-based models demonstrate more comprehensive capabilities and excel in classification, imputation, and anomaly detection tasks. The RNN-based models, while performing well on anomaly detection tasks, show limited effectiveness compared to other model architectures. In contrast, the Transformer-based models have demonstrated highly competitive performance across various time series analysis tasks. This can be attributed to the powerful data modeling capabilities inherent in the transformer architecture, which contribute to its overall and consistently superior performance across diverse time series analysis tasks. It further shows that Transformer-based models hold significant research value and application potential in the field of time series analysis and have emerged as a particularly promising option in the time series domain.
@@
We believe that TSLib can provide useful start code, valuable insights on model properties, and model selection guidance for future research and real-world applications.
Efficiency Analysis We also provide an efficiency comparison to offer deeper insights into the practicality and feasibility of different approaches. Specifically, we measured the training time, inference time, and memory usage of each model in performing long-term forecasting, where all models operated under unified hyperparameter settings. As shown in Figure 10, it is observed that the tokenization strategy also affects the model efficiency, where Autoformer [22] with point-wise tokens is slower than PatchTST [23] with patch-wise tokens and iTransformer [39] with series-wise tokens. Besides, CNN-based models, such as SCINet [130] and TimesNet [3], are much more time-consuming compared to other architectures since they have to adopt diverse kernels to capture intricate temporal patterns. In contrast, MLP-based models and the latest Mamba model stand out for their superior efficiency. Their simpler and more lightweight architectures make them particularly suitable for scenarios with limited computational resources. The above findings underscore the inherent trade-offs between model efficiency and predictive performance, which could also be a valuable guide for selecting models for resource-constrained scenarios.
@@ 6 Future Directions
In this section, we present a discussion on promising directions for time series analysis.
@@ 6.1 Time Series Pre-training
The pretraining-finetuning learning paradigm is a two-stage approach commonly used in nature language processing (NLP) [178], [16], [179], [180], [180] and computer vision (CV) [181], [182], [183], [184]. Pre-training establishes the basis of the abilities of Large Models through unsupervised learning [185, 186]. Fine-tuning can improve the performance of the pre-trained model on a specific task or domain.
@@
Given the scarcity of labeled datasets, self-supervised pre-training [182] has garnered significant attention and has been extensively investigated in the domains of natural language modeling and computer vision. This paradigm significantly reduces labeling expenses and yields considerable benefits for diverse downstream tasks. Notably, recent research has extended self-supervised pre-training methods to time series data, primarily categorizing them into contrastive learning [187] and masked time series modeling [16, 188].
@@
Contrastive learning refers to learning the representations of data by contrasting similar and dissimilar pairs, where similar sample pairs are learned to be close to each other and dissimilar pairs are far apart [187, 189, 190]. Although SimCLR [189] has demonstrated remarkable success in the domain of computer vision, directly applying SimCLR to the field of time series data yields unsatisfactory results due to the insufficient modeling of temporal dependencies. CPC [191] introduces contrastive predictive coding, which utilizes model-predicted features as positive samples and randomly-sampled features as negative samples, to obtain time series representations that are advantageous for downstream tasks. TimeCLR [192] proposes a DTW data augmentation to generate phase shift and amplitude-change phenomena, which can preserve time series structure and feature information. TS-TCC [193] employs efficient data augmentations designed for time-series data, and learns discriminative representations from the proposed Temporal Contrasting module and Contextual Contrasting module. TS2Vec [194] employs a hierarchical contrastive learning method and defines the contrastive loss from both instance-wise and patch-wise perspectives across different augmented context views, resulting in a robust contextual representation for each timestamp. Furthermore. LaST [195] takes the idea of variational inference theory [196] and proposes seasonal-trend representations learning and disentanglement mechanisms. CoST [197] proposes a contrastive learning framework to learn disentangled seasonal-trend representations for long sequence time series data. TF-C [198] develops frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency.
@@
Masked modeling is a reconstruction-based method, which can predict a masked token in a sequence based on the context unmasked part [199, 16, 188]. TST [200] follows the pre-training paradigm proposed in BERT [16] and proposes a pre-training framework for multivariate time series. Further, PatchTST [23] segments the time series data into multiple non-overlapping patches and proposes a patch-level masked modeling approach. SimMTM [201] relates masked modeling to manifold learning and presents a neighborhood aggregation design for reconstruction based on the similarities learned in the series-wise representation space. HiMTM [202] proposes a novel hierarchical masked time series pre-training framework to effectively capture the multi-scale characteristics of time series data. TimeSiam [203] constructs an asymmetric masking reconstruction task to capture intrinsic temporal correlations between randomly sampled past and current subseries and learn internal time-dependent representations based on Siamese networks.
@@ 6.2.1 Time Series Foundation Models
Recent advancements in deep learning, particularly with the emergence of foundation models (FMs), have demonstrated significant progress in natural language processing (NLP) and computer vision (CV) domains [204, 186, 205]. Different from prior deep models, foundation models are pre-trained on massive amounts of data, which enables them to have a wide range of general knowledge learned from diverse domains. Given their success in capturing contextual information and semantic understanding, it is promising to explore a generalized time series foundation model that can effectively learn complex temporal dependencies and capture the underlying dynamics inherent in time series data. Early attempts such as TimeGPT [206], Lag-LlaMa [207], and Timer [157] focus solely on univariate time series data. Nevertheless, in real-world forecasting scenarios, it is crucial to involve additional information that is related to the temporal variation of the target time series and must be taken into account, such as weather conditions or holidays. MOIRAI [208] tries to flatten multivariate time series into a single sequence containing all variates, but its generalization capabilities to other downstream analysis tasks are under-explored. Beyond merely modeling inter-series dependencies, incorporating the relationship between time series and external factors offers a deeper understanding of the data. These external factors can manifest as other modalities, such as textual information or calendar data. Consequently, multimodal learning represents a promising direction for future time series foundation models.
@@ 6.2.2 Adaptation of Large Language Models
Fine-Tuning Pre-trained Language Models Based on the similar sequential nature, fine-tuning pre-trained language models to equip them with time series analysis capabilities has become a promising research topic. When applying LLMs to time series, it is essential to tokenize the time series before feeding it to a pre-trained model. Thus, adapting an LLM for time series included two key components: time series tokenization and efficient fine-tuning methods for time series analysis tasks. LLM4TS [211] proposes a two-stage fine-tuning strategy, including the time-series alignment stage to align LLMs with the nuances of time series data, and the fine-tuning stage for downstream tasks. LLMTime [212] treats time series forecasting as next-token prediction in text and attempts to encode time series data as a string of numerical digits. Recent Chronos [213] introduces a pre-trained probabilistic time series model based on existing Transformer-based language model architectures. Technologically, Chronos tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the model on these tokenized time series via the cross-entropy loss. Benefiting from the generative capability of LLMs, most of the existing research focuses on time series forecasting tasks. GPT4TS [156] proposes a unified framework for diverse time series analysis tasks by using a pre-trained GPT-2 model and fine-tuning the positional embeddings and the parameters of the layer normalization for each analysis task.
Prompting Large Language Models Recent Large Language Models exhibit the strong abilities of in-context learning [147] and instruction following [214]. Therefore, the paradigm of leveraging natural language instructions or task examples to guide the model in addressing novel tasks has emerged as a groundbreaking approach [215, 216, 217], which has become a potential solution for time series analysis tasks [218]. Recent literature, such as PromptCast [218], UniTime [219], and TimeLLM [220] focus on investigating a prompt template to enable LLMs to perform the forecasting task. There are other works represented by Autotimes [53, 221], that attempt to design soft prompts for time series data. However, existing prompting approaches are tailored for forecasting, and how to empower LLMs to perform other time series tasks besides forecasting is relatively unexplored.
Both of the above two directions would require a large-scale training dataset. However, due to the privacy of industrial or IoT scenarios, the accessible time series data is limited, necessitating the exploration of time series generation techniques [222, 213]. Generating high-quality and diverse time series data not only provides valuable resources for model training but also holds the potential to significantly enhance the performance, robustness, and generalization capabilities of these models [223].